OpenStudy (anonymous):

eliminations x + 2y = 9 2x + 6y = -2

jimthompson5910 (jim_thompson5910):

x + 2y = 9 2x + 6y = -2 -2(x + 2y) = -2*9 2x + 6y = -2 -2x-4y = -18 2x + 6y = -2 ------------- 0x + 2y = -20 2y = -20 Solve for y. Then use this value of y to find the solution for x